OpenStudy (anonymous):

w+2x-2y+3z=12 2w+3x+y+2z=10 3w-x+3y+4z=16 -w+4x-2y+4z=20

OpenStudy (anonymous):

what's the question??

OpenStudy (anonymous):

w+2x-2y+3x=12 2w+3x+y+2x=10 3w-x+3y+4z=16 -w+4x-2y+4z=20 my teacher said it should be the same value

OpenStudy (anonymous):

Is this 4 equations 4 unknowns. Solve for w,x,y,z. Also is there supposed to be a z term in the first two?

OpenStudy (anonymous):

yes solve for w,x,y,z and sorry typo 3x and 2x should be 3z and 2z

OpenStudy (anonymous):

What class are you in?

OpenStudy (anonymous):

Linear algebra?

OpenStudy (anonymous):

4th year highschool and i dont know my teacher dont want to tell us she said figure out on ourselves..

OpenStudy (anonymous):

Are you given any values for the 4 variables? If not you basically have to pick one equation and solve in terms of one variable like x, the next in another variable like y=....etc. Once you have your 4 new equations you can start replacing different terms until you have solved for one variable numerically. Or you could set up a matrix.....

OpenStudy (anonymous):

There are a few ways to go about it, not sure what you are expected to do sorry );

OpenStudy (anonymous):

Here's the brute force way. solve w = 12-2x+2y-3z. plug that into the second equation and solve for x. 2(12-2x+2y-3z)+3x+y+2z=10 or 14+5y-4z=x. Now plug w and x into the 3rd equation and solve for y. Then plug w,x,y into the fourth equation and solve for z.

OpenStudy (anonymous):

can u put that in step by step

OpenStudy (anonymous):

but its okay thank u.